Find the area of the region bounded by the graphs of the equations. y = 8x2 + 4, x = 0, x = 2, y = 0Evaluate the definite integral by the limit definition. 7 x dx -6 XEvaluate the definite integral. Use a graphing utility to verify your result. (t1/ dt
Answer #1

Given y=8x²+4 x=0,272,9=0 Alena firaylox 2 ilanteess? t 8 64 3 88 3 y = 3+(302 2} Alea= $% (3+23) dre Boats a les (3x + 39fox de Sondre- (+ux with n=1 I tu حکہ с src- р. ) s froides 13 8 9- 1 (° («s 1) 40 = (35Edt - ff13 at и Si t? ra on • m